OpenStudy (anonymous):

what is 3*f(2)/2 for the function f(x)=5x-2/x^2+3?

OpenStudy (anonymous):

find f(2)

OpenStudy (anonymous):

that would be 8/7 right?

OpenStudy (solomonzelman):

8/7 for the entire thing or just f(2)?

OpenStudy (anonymous):

8/7 for just f(2), but then idk what to do becuase it says to find 3*f(2)/2

OpenStudy (solomonzelman):

you have to divide your result that you got for f(2) (which is 8/7) by 2, and multiply times 2. basically: (8/7) * 3 / 2

OpenStudy (anonymous):

oh ok, thank you ...again:)

OpenStudy (solomonzelman):

just to make sure, please tlel me what you got

OpenStudy (solomonzelman):

(8/7) * 3 / 2 is same as: (8 * 3 / 2 ) /7

OpenStudy (anonymous):

i got 84...is this right?

OpenStudy (solomonzelman):

\(\large\color{black}{ 8 \times 3\div2\div 7 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ 8 \times 3\div2\div 7 }\) \(\large\color{black}{ 4 \times 2\times 3\div2\div 7 }\) \(\large\color{black}{ 4 \underline{~\times 2~}\times 3\underline{~\div2~}\div 7 }\)

OpenStudy (solomonzelman):

\(\large\color{black}{ 4 \times 3\div 7 }\)

OpenStudy (anonymous):

\[\frac{ 8 }{ 7 }\times \frac{ 3 }{ 2 }=?\]